OpenStudy (anonymous):

2 sin u cos u - 2√2 sin u - cos u +√2=0. Find all solutions in the interval {0,360°).

OpenStudy (anonymous):

\[2* \sin u (\cos u -\sqrt{2})- (\cos u -\sqrt{2})=(2\sin u -1)*(\cos u -\sqrt{2})=0\] now: 2sin u - 1=0 or \[\cos u = \sqrt{2}\] it will give you: sin u = 1/2 or

OpenStudy (anonymous):

sin u = 1/2 will give you u = pi/6 and u = 5pi/6 \[\cos u = \sqrt{2}\] will give you u = ... finish it